OpenStudy (anonymous):

find the derivative of f(x)= x/sqrt(x^2 - 4)

OpenStudy (anonymous):

\[(12 x)/(-4+x^2)^(5/2) \]

OpenStudy (anonymous):

∫ xdx/ √( x² -4) Let u = √( x² -4) , u² = x² -4 => 2 udu = 2xdx = ∫ udu / u = ∫ du = u +C = √( x² -4) + c